OpenStudy (anonymous):

Find all absolute and relative extreme values of the function. f(x) = 2x/x - 4 on (-2,2) Choose the correct answer.

OpenStudy (anonymous):

A. The absolute maximum is f(4) = ∞ and the absolute minimum is f(2) = -2. B. The absolute maximum is f(2) = -2 and the absolute minimum is f(-2) = 2/3 . C. The absolute maximum is f(-2) = 2/3 and the absolute minimum is f(2) = -2. D. The absolute maximum is f(-2) = 2/3 and the absolute minimum is f(4) = -∞ .

OpenStudy (abb0t):

Take the first derivative of the function. Use quotient rule.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

8/(x - 4)^2

OpenStudy (anonymous):

so? what can you conclude from that?

OpenStudy (anonymous):

umm idk

OpenStudy (anonymous):

the derivative gives you no critical points, just count the interval end points and compare them.do it

OpenStudy (anonymous):

do you know why the derivative give you no critical point?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

hihi... because it cannot equal 0 at any value of x.

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

you see, there is no x on numerator , how can you set it =0 to compute x?

OpenStudy (anonymous):

ok, the leftover is quite trivial, you can do it, right?

OpenStudy (anonymous):

ok, have fun. bye

OpenStudy (anonymous):

ok